1. a Calculate the force between two charges q1 of

2.0 μC and q2 4.0 μC separated by r = 5.0 cm. (Ans: 29 N)


b Let the force calculated in a be F. In terms of F

and without further calculations, state the force

between these charges when:


i the separation r of the charges is doubled (Ans: F/4)

ii q1 and r are both doubled (F/2)

iii q1, q2 and r are all doubled. (F)


Please explain how these problems were solved.


1
Expert's answer
2021-12-16T17:58:21-0500

(a)

Force

"F=\frac{kq_1q_2}{r^2}"


"F=\frac{9\times10^{9}\times2\times10^{-6}\times4\times10^{-6}}{0.05^2}=28.8N
\\F=28.8N \\ (approximately F=29N)"

(1)

Force

"F=\frac{kq_1q_2}{r_1^2}"

"r_1=2r"

"F=\frac{kq_1q_2}{(2r)^2}=\frac{F}{4}"


(2)

"F=\frac{kq_1q_2}{r_1^2}"

"q_1=2q_1,r=2r"

"F=\frac{2kq_1q_2}{(2r)^2}=\frac{kq_1q_2}{2r^2}=\frac{F}{2}"

(3)

"F=\frac{kq_1q_2}{r_1^2}"

"q_1=2q_1,q_2=2q_2,r=2r"

"F=\frac{k2q_12q_2}{(2r)^2}
\\F=\frac{4kq_1q_2}{4r^2}=F"

"F=\frac{kq_1q_2}{r^2}"